OpenStudy (anonymous):

how do you do 1/10 x 2/5

OpenStudy (anonymous):

whats 1*2

OpenStudy (anonymous):

multiply straight across if that is one tenth times two fifths...

OpenStudy (anonymous):

(1/10) * (2/5) = 1/25 = 0.04

OpenStudy (anonymous):

if you were adding you'd have to find a common denominator.

OpenStudy (anonymous):

Just multiply top to top and bottom to bottom: 2/50=1/25

OpenStudy (anonymous):

wat